7. The equipotential surfaces are drawn around the point charge as shown. (see p (see p. 466) レ, ; (a) Is the charge shown at the center a positive charge or a negative charge? (b) How much work would be needed to push a charge ofa- l џС from B to A?- (c) How much work would be needed to push a charge of +1 μC from B to C? (d) How much energy could be extracted if a charge of +1 HC moves from B to D (0 How much kinetic energy would be gained by a charge of +1 uC if allowed to accelerate downhill from point A to infinity? (e) How much energy could be extracted if a charge of-1 μC moves from B to A? (g) I the equipotential lines are separated by a distance of 1 cm, what is the approximate strength of the clectric field between points A and B? (p. 461) (p. 477) 8. a) Write the capacitance of two parallel plates with area, A, separated by a distance, d, and filled with dielectric constant, K. d. b) What happens to the capacitance if one doubles the area of the plates? c) What happens to the capacitance if one doubles the distance separating the plates? d) What happens to the capacitance if one replaces the air between the plates with glass (K 5)? e) Start with potential energy stored by a capacitor given by U QV and use Q CV to douses halves eliminate Q and write U in terms of C and V f) Start with potential energy stored by the capacitor given by U-% QV and use Q-CV to eliminate V and write U in terms of C and Q # 7.   & if you know the last part of 8 that would be awesome.
Answer #1

7.

(a) Positive

(b)W=q(V_A-V_B)=1cdot (10-7): mu J=3: mu J

(c)

W=q(V_C-V_C)=1cdot (7-7): mu J=0: mu J

(d)

W=q(V_D-V_B)=1cdot (5-7): mu J=-2: mu J

(e)

W=q(V_A-V_B)=-1cdot (10-7): mu J=-3: mu J

(f)

K=(1cdot 10): mu J=10mu J

(g)

10-9 V/m 100V/m d 0.01

8.

(f)

U=rac {1}{2}QV,Q=CV

U=rac {1}{2}Qrac {Q}{C}

U=rac {Q^2}{2C}